### v2

# A Minimal Substitution Basis for the Kalmar Elementary Functions

Abstract
We show that the class of Kalmar elementary functions can be inductively generated from the addition, the integer remainder and the base-two exponentiation, hence improving previous results by Marchenkov and Mazzanti. We also prove that the substitution basis defined by these three operations is minimal. Furthermore, we discuss alternative substitution bases under arity constraints.
Keywords: arithmetic term; closed form; elementary function; elementary operation; explicit formula; Grzegorczyk’s hierarchy.
2020 Mathematics Subject Classification: 03D20 (primary), 03D55, 03B70, 68Q15.
1 Introduction
[…]
Given an integer $r\geq 1$ and a $r$-ary operation $f$ on $\mathbb{N}$, we sometimes write $f$ as $(x_{1},\ldots,x_{r})\mapsto f(x_{1},\ldots,x_{r})$ or even just as $f(x_{1},\ldots,x_{r})$. For example, we sometimes denote the addition as $x+y$.
We also apply the conventions $0^{0}=1$ (see Mendelson [9, Proposition 3.16]) and $\left\lfloor x/0\right\rfloor=0$ (see Mazzanti [8, Section 2.1]), so, in particular, by applying the definition of the integer remainder, $x\bmod y:=x-y\left\lfloor x/y\right\rfloor$, we have that $x\bmod 1=0$ and $x\bmod 0=x$.
Given a set $F$ of operations on $\mathbb{N}$, we denote by $\langle F\rangle$ the closure under substitution of the set $F\cup C\cup P$, where $C$ is the set of constant operations on $\mathbb{N}$ (which can be identified with non-negative integers) and $V$ is the set of projections on $\mathbb{N}$ (which can be identified with variables). And we say that $F$ is a substitution basis for a given set $G$ of operations on $\mathbb{N}$ if, and only if, $\langle F\rangle=G$ (cf. Mazzanti [8, Section 2.2]). For instance, if $S:=\langle x+y,\ x^{y}\rangle$, then $x^{z+2}\in S$: indeed, the projection $(x,y,z)\mapsto z$ and the constant operation $(x,y,z)\mapsto 2$ belong to $S$, so, because of the closure under substitution of $S$, the operation $(x,y,z)\mapsto x^{z+2}$ also belongs to $S$. Observe that some authors do not admit the constant operations as basic elements in a substitution basis (cf. Marchenkov [6, Introduction]), but, as we are already admitting any variable $x_{n}$, it is not unreasonable to also admit any constant $n$.
The Kalmar elementary functions are the operations on $\mathbb{N}$ that can be computed in iterated exponential time (cf. Marchenkov [6, Introduction] and Oitavem [2, Introduction]). They form the class $\mathcal{E}^{3}$ in the Grzegorczyk hierarchy (see Grzegorczyk [1]), which encompasses most of the operations on $\mathbb{N}$ that frequently arise in mathematical contexts. Notable examples include classical number-theoretic functions such as the factorial function $n!$ (see Robinson [3] and Prunescu & Sauras-Altuzarra [11]), the $n$-th prime function $p_{n}$ (see Prunescu & Shunia [15]), Euler’s totient function $\varphi(n)$ (see Prunescu & Sauras-Altuzarra [12]), the greatest common divisor $\gcd(m,n)$ (see Prunescu & Shunia [14]), and the C-recursive integer sequences with rational coefficients (see Prunescu [10] and Prunescu & Sauras-Altuzarra [13]). In practical terms, a significant portion of the real-world computation can be faithfully modeled within the class of Kalmar elementary functions, since the corresponding code avoids unbounded recursion.
For a long time, it was hypothesized that a finite quantity of elementary operations should suffice to form a substitution basis for the class of Kalmar elementary functions. Concretely, it was conjectured that $\langle x+y,\ x\dotdiv y,\ xy,\ \left\lfloor x/y\right\rfloor,\ x^{y}\rangle=\mathcal{E}^{3}$ (see Marchenkov [6, Introduction]). Building upon the foundational contributions of Marchenkov [5], Jones & Matiyasevich [4], and Matiyasevich [7], Mazzanti [8] proved this conjecture. Slightly later, Marchenkov [6, Corollary 2] established that $\langle x+y,\ x\bmod y,\ x^{2},\ 2^{x}\rangle=\mathcal{E}^{3}$. 1 improves Marchenkov’s result by eliminating the squaring operation from the above substitution basis, that is to say, it asserts that $\langle x+y,\ x\bmod y,\ 2^{x}\rangle=\mathcal{E}^{3}$. And Section 3 is devoted to show that this substitution basis is minimal: if $F\subsetneqq\{x+y,x\bmod y,2^{x}\}$, then $\langle F\rangle\subsetneqq\mathcal{E}^{3}$.
In addition, Section 4 analyzes why certain selected sets cannot be substitution bases for the class of Kalmar elementary functions. And, in Section 5, we present some finite substitution bases for the class of univariate Kalmar elementary functions that consist of unary operations only, and a substitution basis for the whole class of Kalmar elementary functions that consists of exactly one binary operation.
2 A substitution basis for the Kalmar elementary functions
We start by stating Theorem 1, which is one of the main results obtained by Marchenkov [6, Corollary 2], already mentioned in Section 1.
Theorem 1.
$\langle x+y,\ x\bmod y,\ x^{2},\ 2^{x}\rangle=\mathcal{E}^{3}$.
Theorem 2 shows how to express squaring in terms of the addition, the integer remainder, and the base-two exponentiation.
Theorem 2.
[…]
Proof.
It is clear that $(2^{x}+x)\mid(2^{x}-x)(2^{x}+x)=2^{2x}-x^{2}$, so $2^{2x}\equiv x^{2}\pmod{2^{x}+x}$ and, as we have that $0\leq x^{2}<2^{x}+x$, we have that
$x^{2}=2^{x+x}\bmod(2^{x}+x).$
Our main result, 1, is now immediate from Theorem 1 and Theorem 2.
Corollary 1.
$\langle x+y,\ x\bmod y,\ 2^{x}\rangle=\mathcal{E}^{3}$.
We conclude this section by providing a proof of 2, which asserts that the truncated subtraction, the product, and the integer division certainly can be expressed in terms of the elements of the base from Theorem 1. This result relies in Theorem 3 and Theorem 4, identities applied by Marchenkov [6] but whose proof appeared to be missing.
Theorem 3.
If $x$ and $y$ are non-negative integers, then
$\left((2^{x+y}+x)\bmod(2^{x+y}+y)\right)\bmod(2^{x+y}+x)=\begin{cases}x-y&\textup{ if }x\geq y,\\ 0&\textup{ if }x<y\end{cases}=x\dotdiv y.$
Proof.
It is clear that
$\begin{cases}\left\{\begin{array}[]{l}0\leq x-y<2^{x+y}+y\\ 2^{x+y}+x=1\cdot(2^{x+y}+y)+(x-y)\end{array}\right.&\textup{ if }x\geq y,\\ \left\{\begin{array}[]{l}0\leq 2^{x+y}+x<2^{x+y}+y\\ 2^{x+y}+x=0\cdot(2^{x+y}+y)+(2^{x+y}+x)\end{array}\right.&\textup{ if }x<y,\end{cases}$
so
$(2^{x+y}+x)\bmod(2^{x+y}+y)=\begin{cases}x-y&\textup{ if }x\geq y,\\ 2^{x+y}+x&\textup{ if }x<y\end{cases}$
and, by again applying that $0\leq 2^{x+y}+x<2^{x+y}+y$ if $x<y$, the statement follows. ∎
Theorem 4.
If $x$ and $y$ are non-negative integers, then
$\left\lfloor x/y\right\rfloor=\left(2(x+1)\left(x\dotdiv(x\bmod y)\right)\right)\bmod(2(x+1)y\dotdiv 1).$
Proof.
If $y=0$, then, by the convention mentioned in Section 1, we have that $\left\lfloor x/y\right\rfloor=\left\lfloor x/0\right\rfloor=0=0\bmod 0=\left(2(x+1)\left(x\dotdiv x\right)\right)\bmod(0\dotdiv 1)=\left(2(x+1)\left(x\dotdiv(x\bmod 0)\right)\right)\bmod(2(x+1)0\dotdiv 1)=\left(2(x+1)\left(x\dotdiv(x\bmod y)\right)\right)\bmod(2(x+1)y\dotdiv 1)$. Otherwise, by applying that $x\dotdiv(x\bmod y)=y\lfloor x/y\rfloor$, we have that $2(x+1)\left(x\dotdiv(x\bmod y)\right)=2xy\lfloor x/y\rfloor+2y\lfloor x/y\rfloor=\lfloor x/y\rfloor(2xy+2y\dotdiv 1)+\lfloor x/y\rfloor$ and thus, by applying that $0\leq\lfloor x/y\rfloor\leq x<2xy+2y\dotdiv 1=2(x+1)y\dotdiv 1$, the statement follows. ∎
Corollary 2.
The operations $x\dotdiv y$, $xy$, and $\left\lfloor x/y\right\rfloor$ belong to $\langle x+y,\ x\bmod y,\ 2^{x}\rangle$.
Proof.
Let $S=\langle x+y,\ x\bmod y,\ 2^{x}\rangle$.
Theorem 4 implies that $\left\lfloor x/y\right\rfloor\in\langle x+y,\ x\dotdiv y,\ x\bmod y,\ 2xy\rangle$.
By applying Theorem 3, we get $x\dotdiv y\in S$.
Thus, by applying Theorem 2, we have that $2xy=(x+y)^{2}\dotdiv(x^{2}+y^{2})\in S$ and hence $\left\lfloor x/y\right\rfloor\in S$.
Finally, notice that $xy=\left\lfloor 2xy/2\right\rfloor\in S$. ∎
3 Minimality of the main substitution basis
In 1, we have seen that the set $\{x+y,x\bmod y,2^{x}\}$ is a substitution basis for the Kalmar elementary functions. In this section we prove that it is, in addition, minimal: $x+y\notin\langle x\bmod y,\ 2^{x}\rangle$, $x\bmod y\notin\langle x+y,\ 2^{x}\rangle$, and $2^{x}\notin\langle x+y,\ x\bmod y\rangle$.
Recall that the definition of substitution basis that we gave in Section 1 permits to use constants when generating functions from a given basis.
Lemma 1.
If $t(x)\in\langle x\bmod y,\ 2^{x}\rangle$, then there is a non-negative integer $B$ such that, for every non-negative integer $a$, the number $t(a)$ is a power of two or $t(a)\leq\max(B,a)$.
Proof.
The proof goes by induction on the elements of $\langle x\bmod y,\ 2^{x}\rangle$.
The statement is obvious for the constants and the variable $x$.
Now, let $t_{1}(x)$ and $t_{2}(x)$ be elements of $\langle x\bmod y,\ 2^{x}\rangle$ that satisfy the property given in the statement: there are two non-negative integers $B_{1}$ and $B_{2}$ such that, for every non-negative integer $a$, the number $t_{1}(a)$ is a power of two or $t_{1}(a)\leq\max(B_{1},a)$, while the number $t_{2}(a)$ is a power of two or $t_{2}(a)\leq\max(B_{2},a)$.
If $t(x)=2^{t_{1}(x)}$, then, for every non-negative integer $a$, the number $t(a)$ is a power of two, and it does not matter how we choose $B$ in this case.
If $t(x)=t_{1}(x)\bmod t_{2}(x)$, then we can take $B=\max(B_{1},B_{2})$: indeed, for every non-negative integer $a$, we have that:
  1. 1.
    if $t_{2}(a)=0$, then $t(a)=t_{1}(a)\bmod 0=t_{1}(a)$ and consequently $t(a)$ is a power of two or $t(a)\leq\max(B_{1},a)\leq\max(B,a)$;
  2. 2.
    if $0<t_{2}(a)\leq\max(B_{2},a)$, then $t(a)<t_{2}(a)\leq\max(B_{2},a)\leq\max(B,a)$;
  3. 3.
    if $t_{1}(a)\leq\max(B_{1},a)$, then $t(a)\leq t_{1}(a)\leq\max(B_{1},a)\leq\max(B,a)$; and
  4. 4.
    if $t_{1}(a)$ and $t_{2}(a)$ are powers of two, then $t(a)$ is a power of two or $t(a)=0\leq\max(B,a)$.
Theorem 5.
$x+y\notin\langle x\bmod y,\ 2^{x}\rangle$.
Proof.
Suppose the contrary. It is then clear that $x+1\in\langle x\bmod y,\ 2^{x}\rangle$. According to Lemma 1, there is some non-negative integer $B$ such that, for every non-negative integer $a$, the number $a+1$ is a power of two or $a+1\leq\max(B,a)$. Therefore, every integer $a>B$ should be a power of two, which is false. ∎
Lemma 2.
If $t(x)\in\langle x+y,\ 2^{x}\rangle$ is non-constant, then $t(x)$ is strictly increasing.
Proof.
The proof goes by induction on the elements of $\langle x+y,\ 2^{x}\rangle$. The statement is obvious for the constants and the variable $x$. And, if $t_{1}(x)$ and $t_{2}(x)$ are strictly increasing elements of $\langle x+y,\ 2^{x}\rangle$, then $t_{1}(x)+t_{2}(x)$ and $2^{t_{1}(x)}$ are also strictly increasing. ∎
Theorem 6.
$x\bmod y\notin\langle x+y,\ 2^{x}\rangle$.
Proof.
Suppose the contrary. Then the operation $x\bmod 2$, which is not strictly increasing, is also a non-constant element of $\langle x+y,\ 2^{x}\rangle$, in contradiction with Lemma 2. ∎
Lemma 3.
If $t(x)\in\langle x+y,\ x\bmod y\rangle$, then $t(x)<Ax+B$ for some non-negative integers $A$ and $B$.
Proof.
The proof goes by induction on the elements of $\langle x+y,\ x\bmod y\rangle$. The statement is obvious for the constants and the variable $x$. And, if $t_{1}(x)$ and $t_{2}(x)$ are elements of $\langle x+y,\ x\bmod y\rangle$ such that $t_{1}(x)<A_{1}x+B_{1}$ and $t_{2}(x)<A_{2}x+B_{2}$ for some non-negative integers $A_{1}$, $A_{2}$, $B_{1}$ and $B_{2}$, then it is clear that $t_{1}(x)+t_{2}(x)<(A_{1}+A_{2})x+(B_{1}+B_{2})$ and $t_{1}(x)\bmod t_{2}(x)<A_{1}x+B_{1}$. ∎
Theorem 7.
$2^{x}\notin\langle x+y,\ x\bmod y\rangle$.
Proof.
Suppose the contrary. Then, Lemma 3 yields the existence of two non-negative integers $A$ and $B$ such that $2^{x}<Ax+B$, which is clearly impossible. ∎
4 Some sets which are not substitution bases
[…]
Lemma 4.
If $t(x)\in\langle 2x,\,x\bmod y,\ 2^{x}\rangle$, then there are two non-negative integers $A$ and $B$ such that, for every non-negative integer $a$, the number $t(a)$ is a power of two or $t(a)\leq A\max(B,a)$.
Proof.
[…]
The statement is obvious for the constants and the variable $x$.
[…]
If $t(x)=2t_{1}(x)$, then, for every non-negative integer $a$, the number $t(a)$ is a power of two or $t(a)\leq 2A_{1}\max(B_{1},a)$, so we can set $A=2A_{1}$ and $B=B_{1}$.
If $t(x)=2^{t_{1}(x)}$, then, for every non-negative integer $a$, the number $t(a)$ is a power of two, and it does not matter how we choose $A$ and $B$ in this case.
If $t(x)=t_{1}(x)\bmod t_{2}(x)$, then we can take $A=\max(A_{1},A_{2})$ and $B=\max(B_{1},B_{2})$: indeed, for every non-negative integer $a$, we have that:
  1. 1.
    if $t_{2}(a)=0$, then $t(a)=t_{1}(a)\bmod 0=t_{1}(a)$ and consequently $t(a)$ is a power of two or $t(a)\leq A_{1}\max(B_{1},a)\leq A\max(B,a)$;
  2. 2.
    if $0<t_{2}(a)\leq A_{2}\max(B_{2},a)$, then $t(a)<t_{2}(a)\leq A_{2}\max(B_{2},a)\leq A\max(B,a)$;
  3. 3.
    if $t_{1}(a)\leq A_{1}\max(B_{1},a)$, then $t(a)\leq t_{1}(a)\leq A_{1}\max(B_{1},a)\leq A\max(B,a)$; and
  4. 4.
    if $t_{1}(a)$ and $t_{2}(a)$ are powers of two, then $t(a)$ is a power of two or $t(a)=0\leq A\max(B,a)$.
Theorem 8.
$\langle 2x,\ x\bmod y,\ 2^{x}\rangle\subsetneqq\mathcal{E}^{3}$.
Proof.
[…]
In a previous version of this article, Theorem 9 was stated as a question, but Emil Jeřábek provided a proof (pers. comm.).
Theorem 9.
$\langle x\bmod y,\ x+1,\ 2^{x}\rangle\subsetneqq\mathcal{E}^{3}$.
Proof.
Let $B$ denote the set of Kalmar elementary functions $f(\vec{x})$ such that, if $r$ is the arity of $f(\vec{x})$, then there exist a non-decreasing computable function $g(x)$ and a number $i\in\{1,\ldots,r\}$ such that $f(\vec{x})\leq g(x_{i})$.
We easily observe that the operations $x+1$, $x\bmod y$, and $2^{x}$ belong to $B$, so, as clearly $\langle B\rangle=B$, we get $\langle x\bmod y,\ x+1,\ 2^{x}\rangle\subseteq B$.
[…]
Corollary 3.
$\langle x\bmod y,\ x\dotdiv y,\ 2^{x}\rangle\subsetneqq\mathcal{E}^{3}$.
Proof.
The proof is the same as in Theorem 9, after replacing $x+1$ with $x\dotdiv y$ in the set. ∎
Note that the proof of Theorem 9 could have been also adapted in order to prove Theorem 5 and Theorem 8.
Finally, we state an open problem, 1.
Problem 1.
Is $\langle x+y,\,\lfloor x/y\rfloor,\,2^{x}\rangle$ equal to $\mathcal{E}^{3}$?
Intuitively, as all three operations displayed in 1 are ultimately monotonic in both arguments, it seems to be challenging to generate periodic behavior (as required for $x\bmod y$) solely from these. But, on the other hand, there are cases in which it is possible: consider, for example, the identity
$2^{x\bmod 2}=\left\lfloor\frac{2^{x}}{2^{\lfloor x/2\rfloor+\lfloor x/2\rfloor}}\right\rfloor,$
which is not hard to prove.
5 Other finite substitution bases with low-arity operations
We conclude the article by discussing the existence of substitution bases for the class of Kalmar elementary functions that consist of only two operations, or even of just one operation.
First of all, we observe that it is not difficult to construct “artificial” substitution bases consisting in just one ternary operation that is not really mathematically relevant, for instance the operation $U(x,y,z)$ defined as
$\begin{cases}2^{x},&\textrm{ if }z=0,\\ x+y,&\textrm{ if }z=1,\\ x\bmod y,&\textrm{ if }z\geq 2,\end{cases}$
as shown in Proposition 1 below. Therefore, we should restrict our attention to unary and binary operations only.
Proposition 1.
$\langle U(x,y,z)\rangle=\mathcal{E}^{3}$.
Proof.
Let $S=\langle U(x,y,z)\rangle$.
Then $2^{x}=U(x,y,0)\in S$, $x+y=U(x,y,1)\in S$, and $x\bmod y=U(x,y,2)\in S$.
Therefore, by applying 1, the statement follows. ∎
Cantor’s pairing function, which we denote by $x\mid y]$, is a bijection that maps each pair $(x,y)$ of non-negative integers into the non-negative integer $(x+y)(x+y+1)/2+x$ (see Prunescu & Sauras-Altuzarra [12, Subsection 4.5]). In addition, we consider in this section the unary operations $l(x)$ and $r(x)$ such that $l([x\mid y])=x$ and $r([x\mid y])=y$.
We now note that a finite set of unary operations cannot be a substitution basis for the class of Kalmar elementary functions, because of the simple reason that a term which is built up from unary function-symbols can contain at most one free variable. But it can be a substitution basis for the class of univariate Kalmar elementary functions, as shown in Theorem 10, which is due to Emil Jeřábek (pers. comm.).
Theorem 10.
There is a finite substitution basis for the class of univariate Kalmar elementary functions consisting of unary operations only.
Proof.
Let $B$ be a finite substitution basis for the class of Kalmar elementary functions consisting of unary and binary operations only (such as that from 1).
Let $d(x)=\left[x\mid x\right]$, $u(x)=\left[l(x)\mid x\right]$, and $v(x)=\left[l(r(x))\mid\left[l(x)\mid r(r(x))\right]\right]$.
Given any two operations $a$ and $b$ in $\langle B\rangle$ such that $a$ is unary and $b$ is binary, let us consider the unary operations $a^{\prime}(x)=\left[a(l(x))\mid r(x)\right]$ and $b^{\prime\prime}(x)=\left[b(l(x),l(r(x)))\mid r(r(x))\right]$.
Let $U$ be the union of the sets $\{l(x),r(x),d(x),u(x),v(x)\}$, $\{a^{\prime}(x):a(x)\textrm{ is a unary operation in }B\}$, and $\{b^{\prime\prime}(x):b(x)\textrm{ is a binary operation in }B\}$.
We claim that $U$ is a finite basis for the univariate Kalmar elementary functions. And it suffices to prove by induction that, for each unary operation $f(x)$ in $\langle B\rangle$, the operation $f^{\prime}(x)$ belongs to $\langle U\rangle$, because then we can express $f(x)$ as $l(f^{\prime}(d(x)))$.
If $f(x)=x$, then $f^{\prime}(x)=[l(x)|r(x)]=x$.
Now, let us consider a unary function $f(x)$ in $\langle U\rangle$ and a unary function $h(x)$ in $B$. In addition, set $F(x)=h(f(x))$. Then $F^{\prime}(x)=(h\circ f)^{\prime}(x)=(h^{\prime}\circ f^{\prime})(x)$, and we have that $f^{\prime}(x)\in\langle U\rangle$ (by the induction hypothesis) and $h^{\prime}(x)\in U$ (by the definition of $U$).
Finally, let us consider two unary functions $f(x)$ and $g(x)$ in $\langle U\rangle$ and one binary function $h(x,y)$ in $B$. In addition, set $F(x)=h(f(x),g(x))$. We observe that
$u(x)=\left[l(x)\mid x\right],$
$(g^{\prime}\circ u)(x)=\left[g(l(x))\mid x\right],$
$(v\circ g^{\prime}\circ u)(x)=\left[l(x)\mid\left[g(l(x))\mid r(x)\right]\,\right],$
$(f^{\prime}\circ v\circ g^{\prime}\circ u)(x)=\left[f(l(x))\mid\left[g(l(x))\mid r(x)\right]\,\right],$
$(h^{\prime\prime}\circ f^{\prime}\circ v\circ g^{\prime}\circ u)(x)=\left[h(f(l(x)),g(l(x)))\mid r(x)\right]=F^{\prime}(x).$
Corollary 4.
There is a substitution basis for the univariate Kalmar elementary functions consisting of exactly eight unary operations.
Proof.
If we restrict the proof of Theorem 10 to the substitution basis $\{x+y,\ x\bmod y,\ 2^{x}\}$ from 1, then the corresponding substitution basis $U$ becomes the union of the sets
$\{l(x),r(x),d(x),u(x),v(x)\},$
$\{a^{\prime}(x):a(x)=2^{x}\}$, and $\{b^{\prime\prime}(x):b(x)\in\{x+y,\ x\bmod y\}\}$, which consists of exactly eight unary operations. ∎
For example, let $f(x)=2x$. Then we can represent the operation $f(x)$ in the basis for the class of univariate Kalmar elementary functions that is given in the proof of Theorem 10 by setting $a(x)=[l(x)+l(r(x))\mid r(r(x))]$: indeed, by performing a few calculations, we obtain $f(x)=l([f(x)\mid x])=l([x+x\mid x])=l(a([x\mid[x\mid x]]))=l(a(u([x\mid x])))=l(a(u(d(x))))$.
Finally, Theorem 11, which is due to Emil Jeřábek (pers. comm.), solves the initial question that was posed at the beginning of this section.
Theorem 11.
There is a basis for the class of Kalmar elementary functions that consists of one binary operation only.
Proof.
By applying Theorem 10, there is an integer $k$ and a substitution basis $B$ for the class of univariate Kalmar elementary functions that consists of exactly $k$ unary operations $f_{0}(x)$, $\ldots$, $f_{k-1}(x)$ such that $f_{0}(x)=2(x+1)$.
In addition, it is easy to prove that, if $i$ is a non-negative integer, then the $i$-th iteration of $f_{0}(x)$, denoted, as usual, as $f_{0}^{i}(x)$, is equal to $2^{i}x+2^{i+1}-2$. For instance, $f_{0}^{0}(x)=x=2^{0}x+2^{0+1}-2$, $f_{0}^{1}(x)=f_{0}(x)=2(x+1)=2^{1}x+2^{1+1}-2$, and $f_{0}^{2}(x)=f_{0}(f_{0}(x))=2(2(x+1)+1)=2^{2}x+2^{2+1}-2$.
Now, let
[…]
which is well-defined because the cases are disjoint: indeed, notice that, if $i$ is a positive integer, then every term of the sequence $f_{0}^{i}(x)$ is even.
We have that $f_{0}(x)=f(x,x)$ and $f_{i}(x)=f(x,f_{0}^{i}(x))$ for every $i\in\{1,\ldots,k-1\}$, so the set $\{f(x,y)\}$ is also substitution basis for the class of univariate Kalmar elementary functions.
Thus, $2x$ and $2x+1$ belong to $\langle f(x,y)\rangle$, whence it follows that $x\mid y]=f(2x,2y+1)\in\langle f(x,y)\rangle$.
Finally, let $s(x)=l(x)+r(x)$ and $m(x)=l(x)\bmod r(x)$.
The operations $s(x)$, $m(x)$, and $2^{x}$ are univariate Kalmar elementary functions, so they belong to $\langle f(x,y)\rangle$ and consequently we have that $x+y=s([x\mid y])\in\langle f(x,y)\rangle$ and $x\bmod y=m([x\mid y])\in\langle f(x,y)\rangle$. Hence, by applying 1, we deduce that $\langle f(x,y)\rangle=\mathcal{E}^{3}$. ∎